OpenStudy (anonymous):

Can someone check this? C = 5/9 ( f - 32 ) 9c = 5(f -32) f - 32 = 9c/5 f = 9c/5 +32 And help with this? Determine how many degrees Fahrenheit 4 degrees Celsius is.

OpenStudy (anonymous):

correct plug in 4 for c

OpenStudy (anonymous):

yes correct

OpenStudy (anonymous):

I dont understand, soo.. 9 * 4 = 5(f-32) f - 32 = 36/5 f= 36/5 + 32 So C would be 39.2? xD

OpenStudy (anonymous):

@roberto5 since NickR got off, is this how I would do the second one?

OpenStudy (anonymous):

one second

OpenStudy (anonymous):

yes correct

OpenStudy (anonymous):

Awesome ! Thanks